OpenStudy (anonymous):

Is there a bijection between the natural numbers and the integers? If there is one, can you please give me an example? I'm drawing a blank.

OpenStudy (jdoe0001):

hmm not sure what you mean all integers are just a subset of the natural set so to that extent, there's a bijection, or one-to-one, if you apply restrictions to the real number set to only non-decimals

OpenStudy (anonymous):

is there an example that takes {1,2,3,4,......} to the integers {....-3,-2,-1,0,1,2,3......} I can't seem to find one that is working.

OpenStudy (jdoe0001):

hmm, not sure... what's being asked

OpenStudy (anonymous):

The question says: "Is there a bijection between the natural numbers N and the integers Z? If so, provide such a bijection. If not, explain carefully why not.

OpenStudy (jdoe0001):

http://www.mathsisfun.com/definitions/natural-number.html <---- I misread you earliers... kinda read real numbers

OpenStudy (jdoe0001):

"is there a bijection between the natural numbers N and the integers Z" meaning the natural numbers, which are really a subset of integers, pretty much just from 0 onwards, so all positive integers can you match them up in a one-to-one relation with the integers set? well, since the natural numbers are a set of all integers and the natural numbers are a subset of all integers, just from 0 onwards then they'd intersect, or biject, from 0 onwards

OpenStudy (jdoe0001):

ahemm.. lemme reword that a bit rather "is there a bijection between the natural numbers N and the integers Z" meaning the natural numbers, which are really a subset of integers, pretty much just from 0 onwards, so all positive integers can you match them up in a one-to-one relation with the integers set? well, since the \(\bf integer\) numbers is a set of all integers and the \(\bf natural\) numbers is a subset of all integers, just from 0 onwards then they'd intersect, or biject, from 0 onwards

OpenStudy (anonymous):

so there is no bijection that can take the natural numbers to the negative integers but there is one that can take it to the positive integers?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

\(\large \begin{array}{ccllll} \mathbb{N}&\mathbb{Z}\\ \hline\\ &-4\\ &-3\\ &-2\\ &-1\\ 0&0\\ 1&1\\ 2&2\\ 3&3\\ 4&4\\ +\infty &+\infty \end{array}\)

OpenStudy (anonymous):

thank you!

OpenStudy (jdoe0001):

yw

OpenStudy (lyrae):

One might think Z should be bigger than N but they actually have the same cardinality. The key to realising this is to write it on a clever form: N: 0 1 2 3 4 5 6 7 8 Z: 0 1 -1 2 -2 3 -3 4 -4 Looking at the relationship between the two rows we can construct the following fuction: \[\phi(n) =\]\[\frac{ -n }{ 2 }\]if n is even. \[\frac{ n + 1 }{ 2 }\]if n is odd. Inserting a \[n \in N\]will always result in a single vaild \[\phi \in Z\]thus there is N for every Z. Also, they are both countably Infinite.

OpenStudy (anonymous):

that will only give negative numbers when n is even

OpenStudy (anonymous):

wait i think i see it

OpenStudy (anonymous):

what about zero? zero isn't a natural number but it is an integer so I need something to give zero

OpenStudy (lyrae):

That's not entirely true. There are different opinions regarding this as N can be defined as the positive integears (0, 1, 2, 3, 4, ..) or the non negative integears (0, 1, 2, 3, 4, ..). The first definition is mostly used in number theorey while the later preferred in set theory and computer science.

OpenStudy (kinggeorge):

Regardless of what the definition is though, you can still construct a bijection. Just take\[\phi(n)=\begin{cases} \frac{n-1}{2},\,n\text{ odd}\\ -\frac{n}{2},\,n\text{ even} \end{cases}\]instead (notice the n-1/2 instead of n+1/2).

OpenStudy (anonymous):

oh gosh thank so much! That helps alot!!!!